# Attention Overlap Is Responsible for The Entity Missing Problem in Text-to-image Diffusion Models!

### v1

@@ Abstract
Keywords Compositional Generation  $\cdot$ Entity Missing  $\cdot$ Diffusion Models
@@ 1 Introduction
In recent years, text-to-image (T2I) diffusion-based (Ho et al., 2020) models such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) and DALL-E (Ramesh et al., 2021, 2022; Betker et al., 2023) have shown promising performance in generating realistic, creative, diverse, and high-quality images from textual descriptions. These models leverage the iterative denoising process, along with text embeddings through the cross-attention mechanism, to generate images used in many real-world applications across various domains, such as controllable image synthesis (Ho and Salimans, 2021; Epstein et al., 2023; Zheng et al., 2023), image editing (Kawar et al., 2023; Zhang et al., 2023), 3D scene production (Poole et al., 2023; Lin et al., 2023; Richardson et al., 2023), video generation (Ho et al., 2022a, b).
However, despite their impressive capabilities, these models often struggle to faithfully capture all the entities, attributes, and relationships described in the input prompt, leading to various compositional misalignments such as entity missing (Chefer et al., 2023; Sueyoshi and Matsubara, 2024; Zhang et al., 2024; Liu et al., 2022; Kim et al., 2023), incorrect attribute binding (Feng et al., 2023; Li et al., 2023; Rassin et al., 2023; Wang et al., 2024), incorrect spatial relationships (Chatterjee et al., 2024; Gokhale et al., 2022; Chen et al., 2024), and numeracy-related problems (Binyamin et al., 2024; Zafar et al., 2024).
Entity missing, in particular, is a prevalent issue where the model fails to faithfully generate one or more entities specified in the textual prompt, significantly hurting the compositional generation performance and limiting the practical usability of T2I models. Previous works have explored both training-based (Ruiz et al., 2022; Wen et al., 2023; Kumari et al., 2022; Zarei et al., 2024) and inference-based Chefer et al. (2023); Sueyoshi and Matsubara (2024); Zhang et al. (2024); Liu et al. (2022); Kim et al. (2023); Karthik et al. (2023); Eyring et al. (2024) approaches to mitigate this problem. While fine-tuning the model on carefully curated datasets can improve compositional alignment, it is computationally expensive, may not generalize well to unseen concepts, and can potentially introduce undesirable properties compared to the primal training phase. On the other hand, inference-time approaches that manipulate the intermediate latent code by optimizing an objective function over cross-attention maps between the input tokens and the image regions have shown promise in addressing entity missing without requiring model retraining. However, there has been a lack of systematic investigation into the optimization objectives that are most effective for this task.
@@
To validate the robustness of our overlap-based metrics, we tested our approach using a diverse set of model backbones containing SD-1.4, SD-2 (Rombach et al., 2022), SD-XL (Podell et al., 2023) and well-established compositional generation benchmarks, such as T2I-CompBench (Huang et al., 2023) and HRS-Bench (Bakr et al., 2023), in addition to the freestyle captions of COCO dataset, all demonstrating the superiority of our overlap-based approach over baselines. Furthermore, manipulating the latent code during the denoising process can lead to undesirable effects, such as reduced image quality and unnatural appearance. To address this issue, we assessed our methods using the Frechet Inception Distance (FID) score (Heusel et al., 2017), Coverage, and Density (Naeem et al., 2020), all demonstrating only a slight degradation in quality compared to the best-performing baseline.
@@ 2.1 Visual Compositional Generation
Compositionality, the study of how basic components combine to form entirely new concepts (Peters et al., 2017), is a fundamental characteristic of the natural world and a crucial aspect of human cognition, allowing individuals to reason and adapt to novel situations (Phillips and Wilson, 2010; Frankland and Greene, 2020; Reverberi et al., 2012). Building on this concept, compositional generation has emerged as a significant area of research in recent years, particularly in developing T2I models. Numerous studies have focused on advancing and evaluating T2I models to improve and assess their ability to handle compositional generation (Huang et al., 2023; Bakr et al., 2023; Li et al., 2024). These studies indicate that despite the impressive capabilities of T2I models in generating realistic and high-quality images, they sometimes fail to produce outputs that fully align with the provided textual input. This misalignment is particularly evident when the model encounters complex or novel compositions poorly represented in the training data, leading to compositional generation failures such as entity missing, incorrect attribute binding, incorrect spatial relationships, and numeracy-related issues.
Entity missing occurs when at least one entity specified in the textual prompt is not completely represented in the generated image (Chefer et al., 2023; Sueyoshi and Matsubara, 2024; Zhang et al., 2024; Liu et al., 2022; Kim et al., 2023). For example, considering the prompt “A cat sitting on a chair next to a dog”, the generated image may only contain the chair, with the cat either missing entirely or only partially visible. This problem typically arises when the model does not allocate sufficient attention to certain entities during the generation process, leading to their partial or total absence in the final output. On the other hand, incorrect attribute binding refers to the mis-association of an attribute, such as color, texture, shape, and size, with the corresponding entity (Feng et al., 2023; Li et al., 2023; Rassin et al., 2023; Wang et al., 2024; Wu et al., 2023; Park et al., 2024). For example, in response to the prompt “A green apple and a red ball” the T2I model might generate an image where the apple is red and the ball is green, incorrectly swapping the attributes. This problem often stems from improper alignment in the attention maps of the entity and its corresponding attribute during the generation process.
@@ 2.2 Training-Free Methods for Compositional Generation
Training-free methods provide a promising, on-the-fly solution for addressing compositional generation failures, such as missing entities, in T2I diffusion models. These approaches eliminate the need for costly fine-tuning while avoiding its potential drawbacks while preserving the generation capability of the original model. Some of these methods leverage the multiple generation paradigm (Karthik et al., 2023; Liu et al., 2024), allowing the T2I model to produce several parallel latent codes during the denoising process. The optimal result is then selected based on an evaluation metric applied either to the final image or to intermediate latent embeddings. However, generating multiple instances significantly increases inference time and makes these methods resource-intensive.
@@
For example, the Attend-and-Excite method (Chefer et al., 2023) employs an intensity-based loss function designed to prevent the weakening of the entity with the lowest maximum attention score, while Divide-and-Bind (Li et al., 2023) uses a Kullback–Leibler (KL) divergence loss function to ensure the proper association of attributes with corresponding entities. Additionally, Predicated Diffusion (Sueyoshi and Matsubara, 2024) interprets the intended meaning of the input prompt as propositions based on predicate logic, leading to a differentiable loss function that guides the image generation process to better satisfy the propositions. Lastly, Attention Regulation (Zhang et al., 2024) formulates a constrained optimization problem to minimize attention mismatches, aligning attention maps more closely with the input text prompt.
@@ 5 Method
The investigation presented in Section 4.4 aimed to clarify the relationship between various metrics applied to cross-attention maps and the occurrence of missing entities. The analysis revealed compelling evidence highlighting the importance of overlap-based metrics, such as IoU, $D_{CoM}$, $D_{KL}$, and CC, in reliably assessing performance related to mitigating missing entities. Building on insights from prior studies (Chefer et al., 2023; Li et al., 2023; Sueyoshi and Matsubara, 2024; Zhang et al., 2024) and leveraging the advantages of training-free methods, we propose a loss function-based approach. This method minimizes a loss function during denoising steps to reduce the overlap between the attention maps of different entities, thereby improving the model’s ability to faithfully represent all entities mentioned in the input text prompt.
@@ 6.1 Experimental Setup
We compared our method to several existing training-free methods, including state-of-the-art ones for improving compositional generation in T2I diffusion models, including Attend-and-Excite (Chefer et al., 2023), Divide-and-Bind (Li et al., 2023), Structured Diffusion (Feng et al., 2023), Predicated Diffusion (Sueyoshi and Matsubara, 2024), Attention-Reg (Zhang et al., 2024), and Composable Diffusion Liu et al. (2022). Moreover, we evaluated our method against a well-known fine-tuning-based approach, Dense Diffusion (Kim et al., 2023), which incorporates user-defined layouts along with the textual input. It is important to note that our proposed overlap-based approach requires no additional input and functions as a plug-in module for any diffusion-based backbone. For all baseline methods, we utilized the official implementations provided by the authors, following their recommended hyperparameter settings. Further details about each baseline are provided in Section 2.2 and Appendix C.1.
@@ 6.3 Qualitative Results
In contrast, our proposed overlap-based methods consistently produce images that accurately capture the composition outlined in the input prompts. By minimizing the overlap between attention maps of distinct entities, our approach ensures that each subject is well-separated within the attention map space and correctly associated with its corresponding attributes. Further qualitative results using additional backbones (Stable Diffusion XL, Stable Diffusion v2, Stable Diffusion v2-base are provided in Appendix E.
@@ 7 Conclusion
Through this work, we conducted a comprehensive study of the entity missing problem, a critical compositional generation failure mode in T2I diffusion models. By investigating three hypothetical sources of the problem through the lens of cross-attention mechanisms - attention intensity, attention spread, and attention overlap - we identified the crucial role that overlap between attention maps of entities play in the entity missing problem. Building upon this insight, we proposed four overlap-based loss functions: $IoU$, $D_{CoM}$, $D_{KL}$, and $CC$, which can be applied over the latent embedding as training-free approaches to mitigate the entity missing issue during denoising process. Extensive experiments on a diverse set of prompts from our proposed dataset COCO-Comp to free-style COCO captions and well-known compositional benchmarks T2I-Compbench and HRS-Bench, which all together span various linguistic structures and entity counts, demonstrated that our proposed methods significantly outperform state-of-the-art baselines in faithfully generating entities, as measured by commonly-used evaluation metrics such as human evaluation, visual question answering (VQA), image captioning score, and CLIP similarity. These results suggest that attention overlap is a primary factor contributing to the entity missing problem in T2I diffusion models. We also evaluate the quality and naturalness of the images generated by our overlap-based methods using the FID score, Coverage, and Density metrics. All metrics indicate that the degradation in quality and diversity is negligible. Furthermore, qualitative results suggest that the application of overlap-based loss functions during the denoising steps does not introduce undesirable biases, such as unnecessary and unnatural spatial distances between entities. While there are limitations to the proposed method (Appendix G), the proposed overlap-based loss functions provide an effective and flexible plug-and-play framework for addressing the entity missing problem without the need for expensive fine-tuning or retraining of the model. We believe this work paves the way for further research into training-free approaches for enhancing the compositional generation capabilities of T2I diffusion models, ultimately leading to more reliable and semantically coherent image synthesis.
@@ Appendix A COCO-Comp Dataset
Unfortunately, many previous works have utilized datasets with limited diversity regarding the entities and categories represented. In response, we emphasize the importance of covering a diverse set of entities spanning a wide range of categories, including animals, foods, furniture, and more, by adopting 71 objects from the COCO dataset Lin et al. [2014] as our entity set. These diverse entities, detailed in table 7, can comprehensively assess the ability of the T2I model to encounter the complexity of real-world environments and even unseen combinations of entities, which is neglected in the previous works. Notably, we filtered out 20 entities from the 91 entities represented in the COCO dataset because they were two-part words, such as “wine glass” or plural, such as “scissors”.

### v2

@@ 1 Introduction
In recent years, text-to-image (T2I) diffusion-based (Ho et al., 2020) models such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) and DALL-E (Ramesh et al., 2021; 2022; Betker et al., 2023) have shown promising performance in generating realistic, creative, diverse, and high-quality images from textual descriptions. These models leverage the iterative denoising process, along with text embeddings through the cross-attention mechanism, to generate images used in many real-world applications across various domains, such as controllable image synthesis (Ho & Salimans, 2021; Epstein et al., 2023; Zheng et al., 2023), image editing (Kawar et al., 2023; Zhang et al., 2023), 3D scene production (Poole et al., 2023; Lin et al., 2023; Richardson et al., 2023), video generation (Ho et al., 2022a; b).
However, despite their impressive capabilities, these models often struggle to faithfully capture all the entities, attributes, and relationships described in the input prompt, leading to various compositional misalignments such as entity missing (Chefer et al., 2023; Sueyoshi & Matsubara, 2024; Zhang et al., 2024; Liu et al., 2022; Kim et al., 2023), incorrect attribute binding (Feng et al., 2023; Li et al., 2023; Rassin et al., 2023; Wang et al., 2024), incorrect spatial relationships (Chatterjee et al., 2024; Gokhale et al., 2022; Chen et al., 2024), and numeracy-related problems (Binyamin et al., 2024; Zafar et al., 2024).
Entity missing, in particular, is a prevalent issue where the model fails to faithfully generate one or more entities specified in the textual prompt, significantly hurting the compositional generation performance and limiting the practical usability of T2I models. Previous works have explored both training-based (Ruiz et al., 2022; Wen et al., 2023; Kumari et al., 2022; Zarei et al., 2024) and inference-based Chefer et al. (2023); Sueyoshi & Matsubara (2024); Zhang et al. (2024); Liu et al. (2022); Kim et al. (2023); Karthik et al. (2023); Eyring et al. (2024) approaches to mitigate this problem. While fine-tuning the model on carefully curated datasets can improve compositional alignment, it is computationally expensive, may not generalize well to unseen concepts, and can potentially introduce undesirable properties compared to the primal training phase. On the other hand, inference-time approaches that manipulate the intermediate latent code by optimizing an objective function over cross-attention maps between the input tokens and the image regions have shown promise in addressing entity missing without requiring model retraining. However, there has been a lack of systematic investigation into the optimization objectives that are most effective for this task.
@@
To validate the robustness of our overlap-based metrics, we tested our approach using a diverse set of model backbones containing SD-1.4, SD-2 (Rombach et al., 2022), SD-XL (Podell et al., 2023) and well-established compositional generation benchmarks, such as T2I-CompBench (Huang et al., 2023) and HRS-Bench (Bakr et al., 2023), in addition to the freestyle captions of COCO dataset, all demonstrating the superiority of our overlap-based approach over baselines. Furthermore, manipulating the latent code during the denoising process can lead to undesirable effects, such as reduced image quality and unnatural appearance. To address this issue, we assessed our methods using the Frechet Inception Distance (FID) score (Heusel et al., 2017), Coverage, and Density (Naeem et al., 2020), all demonstrating only a slight degradation in quality compared to the best-performing baseline. Finally, we evaluated the effectiveness of the proposed overlap-based methods in addressing other challenges in visual compositional generation, including incorrect attribute binding, incorrect spatial relationship, and numeracy-related issues. The results demonstrated that solely improving entity preservation can effectively contribute to mitigating these compositional generation challenges.
@@ 2.1 Visual Compositional Generation
Compositionality, the study of how basic components combine to form entirely new concepts (Peters et al., 2017), is a fundamental characteristic of the natural world and a crucial aspect of human cognition, allowing individuals to reason and adapt to novel situations (Phillips & Wilson, 2010; Frankland & Greene, 2020; Reverberi et al., 2012). Building on this concept, compositional generation has emerged as a significant area of research in recent years, particularly in developing T2I models. Numerous studies have focused on advancing and evaluating T2I models to improve and assess their ability to handle compositional generation (Huang et al., 2023; Bakr et al., 2023; Li et al., 2024). These studies indicate that despite the impressive capabilities of T2I models in generating realistic and high-quality images, they sometimes fail to produce outputs that fully align with the provided textual input. This misalignment is particularly evident when the model encounters complex or novel compositions poorly represented in the training data, leading to compositional generation failures such as entity missing, incorrect attribute binding, incorrect spatial relationships, and numeracy-related issues.
Entity missing occurs when at least one entity specified in the textual prompt is not completely represented in the generated image (Chefer et al., 2023; Sueyoshi & Matsubara, 2024; Zhang et al., 2024; Liu et al., 2022; Kim et al., 2023). For example, considering the prompt “A cat sitting on a chair next to a dog”, the generated image may only contain the chair, with the cat either missing entirely or only partially visible. This problem typically arises when the model does not allocate sufficient attention to certain entities during the generation process, leading to their partial or total absence in the final output (Chefer et al., 2023). On the other hand, incorrect attribute binding refers to the mis-association of an attribute, such as color, texture, shape, and size, with the corresponding entity (Feng et al., 2023; Li et al., 2023; Rassin et al., 2023; Wang et al., 2024; Wu et al., 2023; Park et al., 2024). For example, in response to the prompt “A green apple and a red ball” the T2I model might generate an image where the apple is red and the ball is green, incorrectly swapping the attributes. This problem often stems from improper alignment in the attention maps of the entity and its corresponding attribute during the generation process.
@@ 2.2 Training-Free Methods for Compositional Generation
While several fine-tuning-based approaches (Sun et al., 2023; Jiang et al., 2024; Zarei et al., 2024; Guo et al., 2024; Wen et al., 2023) have been proposed to mitigate failure modes in visual compositional generation, these methods are often resource-intensive, requiring the careful curation of appropriate datasets. Additionally, they may introduce new challenges, such as overfitting to the fine-tuning data and the potential loss of beneficial properties inherent in the T2I model from its initial training phase.
On the other hand, training-free methods provide a promising, on-the-fly solution for addressing compositional generation failures, such as missing entities, in T2I diffusion models. These approaches eliminate the need for costly fine-tuning while avoiding its potential drawbacks while preserving the generation capability of the original model. Some of these methods leverage the multiple generation paradigm (Karthik et al., 2023; Liu et al., 2024), allowing the T2I model to produce several parallel latent codes during the denoising process. The optimal result is then selected based on an evaluation metric applied either to the final image or to intermediate latent embeddings. However, generating multiple instances significantly increases inference time and makes these methods resource-intensive.
@@
For example, the Attend-and-Excite method (Chefer et al., 2023) employs an intensity-based loss function designed to prevent the weakening of the entity with the lowest maximum attention score, while Divide-and-Bind (Li et al., 2023) uses a Kullback–Leibler (KL) divergence loss function to ensure the proper association of attributes with corresponding entities. Additionally, Predicated Diffusion (Sueyoshi & Matsubara, 2024) interprets the intended meaning of the input prompt as propositions based on predicate logic, leading to a differentiable loss function that guides the image generation process to better satisfy the propositions. Lastly, Attention Regulation (Zhang et al., 2024) formulates a constrained optimization problem to minimize attention mismatches, aligning attention maps more closely with the input text prompt.
@@ 5 Method
The investigation presented in Section 4.4 aimed to clarify the relationship between various metrics applied to cross-attention maps and the occurrence of missing entities. The analysis revealed compelling evidence highlighting the importance of overlap-based metrics, such as IoU, $D_{CoM}$, $D_{KL}$, and CC, in reliably assessing performance related to mitigating missing entities. Building on insights from prior studies (Chefer et al., 2023; Li et al., 2023; Sueyoshi & Matsubara, 2024; Zhang et al., 2024) and leveraging the advantages of training-free methods, we propose a loss function-based approach. This method minimizes a loss function during denoising steps to reduce the overlap between the attention maps of different entities, thereby improving the model’s ability to faithfully represent all entities mentioned in the input text prompt.
@@ 6.1 Experimental Setup
We compared our method to several existing training-free methods, including state-of-the-art ones for improving compositional generation in T2I diffusion models, including Attend-and-Excite (Chefer et al., 2023), Divide-and-Bind (Li et al., 2023), Structured Diffusion (Feng et al., 2023), Predicated Diffusion (Sueyoshi & Matsubara, 2024), Attention-Reg (Zhang et al., 2024), and Composable Diffusion Liu et al. (2022). Moreover, we evaluated our method against a well-known fine-tuning-based approach, Dense Diffusion (Kim et al., 2023), which incorporates user-defined layouts along with the textual input. It is important to note that our proposed overlap-based approach requires no additional input and functions as a plug-in module for any diffusion-based backbone. For all baseline methods, we utilized the official implementations provided by the authors, following their recommended hyperparameter settings. Further details about each baseline are provided in Section 2.2 and Appendix C.1.
@@ 6.2 Quantitative Results
Background Entites.
While well-known benchmarks for visual compositional generation typically exclude background entities, we extended the scope of evaluation by conducting an experiment incorporating background elements. Specifically, we utilized the CoM distance as the loss function over 96 prompts structured in the format: “a {foreground entity} and a {background entity}”. Stable Diffusion XL served as the backbone model, the foreground entity was chosen from the COCO dataset, and the background entity was varied across six categories: ‘sea’, ‘beach’, ‘desert’, ‘jungle’, ‘sky’, and ‘mountain’. Table 6 presents the results from human assessments, revealing over $4\%$ improvement in both harsh and soft evaluation scenarios.
Table 6: Quantitative comparison between the overlap-based method, CoM distance, and the state-of-the-art T2I model, Stable Diffusion XL, over prompts containing background entities such as ‘sea’, ‘beach’, ‘desert’, ‘jungle’, ‘sky’, and ‘mountain’ using soft and harsh human judgment.
Method Human Score (S) ($\uparrow$) Human Score (H) ($\uparrow$)
Stable Diffusion 91.7% 89.6%
CoM Distance 96.1% 94.8%
@@
Other Visual Compositional Generation Failure Modes.
We further evaluated the effectiveness of the proposed overlap-based metrics in addressing other compositional generation challenges, including incorrect attribute binding, incorrect spatial relationship, and numeracy-related issues. For attribute binding and spatial relationship tasks, we used 100 randomly selected prompts from the color and 2D-spatial sections of the T2I-CompBench benchmark, respectively. To generate numeracy-related images, we utilized 100 prompts in the format “number entity and number entity”, where the number tokens were one of ‘one’, ‘two’, ‘three’, or ‘four’, and the entity tokens were selected from the COCO dataset. The human assessment-based results presented in Table 8 were obtained using the CoM distance metric as the loss function, applied exclusively to entity-related attention maps, with Stable Diffusion v2 serving as the backbone. Notably, for numeracy-related problems, the mean absolute error (MAE) was calculated as the absolute difference between the ground truth number in the prompt and the number of generated entities in the image across all prompts. As expected, entity missing is a fundamental issue in visual compositional generation, and solely enhancing entity preservation can significantly contribute to mitigating other related challenges.
Table 8: Quantitative comparison between the overlap-based method, CoM distance, and the state-of-the-art T2I model, Stable Diffusion v2, over other visual compositional generation tasks: attribute binding, spatial relationship, and numeracy. All the results are based on harsh human assessments. For the attribute binding and spatial relationship tasks, the success rate of the method in accurately generating the corresponding attribute or relation is evaluated. In contrast, for the numeracy task, the mean absolute error (MAE) between the number of entities generated in the images and the ground truth number specified in the prompts is assessed.
Method Attribute Binding ($\uparrow$) Spatial Relationship ($\uparrow$) Numeracy ($\downarrow$)
Stable Diffusion 59.75% 56.33% 2.74
CoM Distance 73.75% 63.66% 2.09
@@ 6.3 Qualitative Results
In contrast, our proposed overlap-based methods consistently produce images that accurately capture the composition outlined in the input prompts. By minimizing the overlap between attention maps of distinct entities, our approach ensures that each subject is well-separated within the attention map space and correctly associated with its corresponding attributes. Further qualitative results using additional backbones (Stable Diffusion XL, Stable Diffusion v2, Stable Diffusion 1.4, Stable Diffusion v2-base) and datasets (COCO-Comp, COCO, and the prompts containing background entities) are provided in Appendix E.
@@ 7 Conclusion
Through this work, we conducted a comprehensive study of the entity missing problem, a critical compositional generation failure mode in T2I diffusion models. By investigating three hypothetical sources of the problem through the lens of cross-attention mechanisms - attention intensity, attention spread, and attention overlap - we identified the crucial role that overlap between attention maps of entities play in the entity missing problem. Building upon this insight, we proposed four overlap-based loss functions: $IoU$, $D_{CoM}$, $D_{KL}$, and $CC$, which can be applied over the latent embedding as training-free approaches to mitigate the entity missing issue during denoising process. Extensive experiments on a diverse set of prompts from our proposed dataset COCO-Comp to free-style COCO captions and well-known compositional benchmarks T2I-Compbench and HRS-Bench, which all together span various linguistic structures and entity counts, demonstrated that our proposed methods significantly outperform state-of-the-art baselines in faithfully generating entities, as measured by commonly-used evaluation metrics such as human evaluation, visual question answering (VQA), image captioning score, and CLIP similarity. These results suggest that attention overlap is a primary factor contributing to the entity missing problem in T2I diffusion models. We also evaluate the quality and naturalness of the images generated by our overlap-based methods using the FID score, Coverage, and Density metrics. All metrics indicate that the degradation in quality and diversity is negligible. Furthermore, qualitative results suggest that the application of overlap-based loss functions during the denoising steps does not introduce undesirable biases, such as unnecessary and unnatural spatial distances between entities. While there are limitations to the proposed method (Appendix G), the proposed overlap-based loss functions provide an effective and flexible plug-and-play framework for addressing the entity missing problem without the need for expensive fine-tuning or retraining of the model. Furthermore, our experiments indicated that the exclusive enhancement of entity preservation can contribute to the alleviation of other challenges in compositional generation, such as incorrect attribute binding, incorrect spatial relationship, and issues related to numeracy. We believe this work paves the way for further research into training-free approaches for enhancing the compositional generation capabilities of T2I diffusion models, ultimately leading to more reliable and semantically coherent image synthesis.
@@ Appendix A COCO-Comp Dataset
Unfortunately, many previous works have utilized datasets with limited diversity regarding the entities and categories represented. In response, we emphasize the importance of covering a diverse set of entities spanning a wide range of categories, including animals, foods, furniture, and more, by adopting 71 objects from the COCO dataset Lin et al. (2014) as our entity set. These diverse entities, detailed in table 9, can comprehensively assess the ability of the T2I model to encounter the complexity of real-world environments and even unseen combinations of entities, which is neglected in the previous works. Notably, we filtered out 20 entities from the 91 entities represented in the COCO dataset because they were two-part words, such as “wine glass” or plural, such as “scissors”.